OpenStudy (anonymous):

Your Math class has 32 students. You would like to treat the class to cookies. You would like each person to have 5 cookies. Which expression would be a quick way to figure out how many cookies you would need to bake? 5(32 + 2) 5(30 + 2) 5(32 − 2) 5(30 − 2)

OpenStudy (phi):

if it were 1 student the answer is 5 cookies 2 then 5+5 =10 3 then 5+5+5= 15 4 then 5+5+5+5= 20 it looks like 5 times the number of students. so 5*32 none of the answers look like that... but let's add the numbers inside the parentheses: 5(32+2)= 5*34 that is not right. Try the rest of them...